OpenStudy (luigi0210):

Integration:

OpenStudy (luigi0210):

\[\LARGE \int~x^5\sqrt{x^2+4}~dx\]

OpenStudy (anonymous):

sub x = 2tanu dx = 2sec^2u du

OpenStudy (anonymous):

\[4 \int\limits 32\tan^5u \sec^3u du ~~ = 128 \int\limits \tan^5u \sec^3u du\]

OpenStudy (anonymous):

Use the trig identitiy tan^2x = sec^2x -1 \[128 \int\limits tanu \sec^3u (\sec^2u - 1)^2 du\] sub t = sec u dt = tanusecu du \[= 128 \int\limits t^2(t^2-1)^2 dt\] Expand it \[= 128 \int\limits (t^6-2t^4+t^2) dt\] Pretty straight forward from there, and ofc you'd leave >_>

OpenStudy (luigi0210):

I'm right here, I never left bro. And that sub the only way? Or is it just the "best" sub?

OpenStudy (anonymous):

There's always more than one way, but I wanted to do it that way, it's good practice. And omg you changed your picture, I usually expect a baby luigi or something, quit changing your pic.

OpenStudy (anonymous):

You could try parts, not sure how far you'll get but it's worth a shot.

OpenStudy (luigi0210):

Thanks bud :)